OpenStudy (anonymous):

this equation is not Bernoulli's Equation. Since, there exists " 1 " in the right side

OpenStudy (amistre64):

this would prolly be more effective within the original post ... as is it just comes across as a crazy persons rant :)

OpenStudy (agreene):

It sounds like you're saying this--or something. \[\large C_p=\frac{p-p_e}{\frac12 pV^2_e}+1\]

OpenStudy (anonymous):

Please, solve the following differential equation: y' + p(x)y=q(x) \sqrt{y}\ - 1